# Generative Hierarchical Temporal Transformer for Hand Action Recognition and Motion Prediction

### v2

@@ Abstract
We present a novel framework that concurrently tackles hand action recognition and 3D future hand motion prediction. While previous works focus on either recognition or prediction, we propose a generative Transformer VAE architecture to jointly capture both aspects, facilitating realistic motion prediction by leveraging the short-term hand motion and long-term action consistency observed across timestamps. To ensure faithful representation of the semantic dependency and different temporal granularity of hand pose and action, our framework is decomposed into two cascaded VAE blocks. The lower pose block models short-span poses, while the upper action block models long-span action. These are connected by a mid-level feature that represents sub-second series of hand poses. Our framework is trained across multiple datasets, where pose and action blocks are trained separately to fully utilize pose-action annotations of different qualities. Evaluations show that on multiple datasets, the joint modeling of recognition and prediction improves over separate solutions, and the semantic and temporal hierarchy enables long-term pose and action modeling.
@@ 1 Introduction
Understanding dynamic hand poses and actions is a fundamental task for human-robot interaction and VR/AR applications. While there is tremendous progress in recognizing action and predicting future hand motion as a sequence of frame-wise poses, most of the literature tackle only either the recognition [13, 43, 52, 53] or prediction [2, 26, 25] task. On the other hand, recent works demonstrate the benefits of bridging recognition and prediction. For example, [39, 44] learn motion priors through future pose prediction, and apply them to refining pose estimated in videos. Other works [36, 37, 48, 28] learn generative models of action-conditioned motion generation, which can be used to synthesize data for recognition training. Nevertheless, the separate solutions for recognition and prediction tend to overfit to specific data distributions, lacking robust performance generalization (Sec. 4.3).
@@
Figure 1: Jointly modeling recognition and prediction, while following the semantic dependency and temporal granularity for hand pose-action. For recognition, (1$\to$(2) moves up from short clips to long spans for input pose refinement and action recognition respectively. For motion prediction, two paths are available: (1$\to$(4) leverages short-term motion regularity, and (1$\to$(2$\to$(3$\to$(4) enables long-term action guided prediction.
We present a framework with generative Transformer VAE architecture to jointly model both action recognition and motion prediction. The joint modeling enables exploiting the temporal regularity of short-term hand motion and long-term action shared through the observation (e.g. hand poses to reach into the can, indicating the action of taking out a chip) and future timestamps (e.g. predicted poses to grab and pull out a chip, thus completing the action, see Fig. 1). Therefore, the model can address various tasks including input pose refinement, action recognition, and future motion prediction. Additionally, it can enhance the performance on these tasks compared to individual models (Sec. 4.3). Specifically, the transformer encoder is responsible for producing outputs for recognition, while the decoder is responsible for prediction. The VAE latent space connects the two by extracting regular motion and action priors, which are shared between the observed sequences and future motion.
@@
Moreover, when it comes to pose-action modeling, extensive previous works [52, 8, 53, 47, 25, 14, 36, 37, 48, 28] demonstrate the importance of capturing the semantic dependency of hand sequences in different temporal granularities, i.e. the instantaneous pose and the action over seconds. For example, [52, 8, 53, 47] first locate the hand motion (e.g. hand poses to reach into the can, grab a chip, and pull out the chip) from the observed video, and use it to recognize the performed action (e.g. take out a chip). [52] further proposes Hierarchical Temporal Transformer to model the semantic dependency and temporal granularity of pose and action with two cascaded encoders respectively working on short and long time spans, for effective hand pose estimation and action recognition. On the generation side, [14, 36, 37, 48, 28] generate more realistic motion by conditioning on action.
@@
Inspired by [52], in addition to jointly modeling both recognition and prediction, we faithfully respect the semantic dependency and temporal granularity of hand pose-action, which improves over flattened models (Sec. 4.4). Our framework, named Generative Hierarchical Temporal Transformer (G-HTT), consists of two cascaded blocks that have the same generative transformer VAE structure but focus on different semantic and temporal granularities: the lower pose block models hand poses over short time spans, and the upper action block models action over long time spans. Meanwhile, the two blocks are bridged by a middle-level representation, i.e. the pose block VAE latent code, which encodes clip-wise motion over a subsecond span (Fig. 1). The hierarchy further brings training flexibility, as we can train the blocks separately, which not only reduces training computational demands but also allows for using training data with pose-action annotations of different qualities (Secs. 3.4 and 4.5).
@@
We train and evaluate the framework across different datasets of two-hand interactions, including H2O [22] for daily activities, Assembly101 [42] and AssemblyHands [35] for (dis-)assembling take-part toys. At test time, given a 3D hand pose sequence (e.g. per-frame estimations from the observed RGB video), we first refine it by leveraging the short-term hand motion regularity, (Fig. 1, (1)). Next, we aggregate the clip-wise motions for action recognition (Fig. 1, (1$\to$(2)). Finally, we decode the observed action and motions into a sequence of future middle-level features for motion prediction (Fig. 1, (1$\to$(2$\to$(3$\to$(4)). Evaluation results across datasets show that our framework can solve recognition problems from various camera views, and produce plausible future hand poses over time. The contribution of this paper can be summarized as follows:
@@
  • A joint modeling of hand action recognition and motion prediction with generative transformer VAEs, which exploits the temporal regularity of motion and action shared between the observed and future timestamps, thus contributing to robust performance generalization,
  • A specialized architecture composed of two generative blocks that model the semantic dependency and temporal granularity of pose-action, thus improving performance and training flexibility, and
@@
  • A comprehensive evaluation of the system on tasks such as 3D hand pose refinement, action recognition, and 3D motion prediction, validating the performance and design of our framework.
@@ 2 Related Works
Action recognition and 3D Hand pose estimation
@@
Massive literature addresses recognition from the visual observation, with tasks at different semantic levels. For example, a series of works aim to recover the 3D hand skeleton or mesh from the visual input, where the spatial correlation within a single-frame is well exploited [58, 19, 46, 23, 33, 54], and the motion cue along the temporal dimension is further leveraged for improved robustness under occlusion and truncation  [34, 3, 10, 15, 51, 16]. Meanwhile,  [5, 12, 13, 11, 43] focus on a higher semantic level, where they extract the spatial-temporal feature from the input frames to recognize the semantic hand or body action.
@@
Moreover, many notice and exploit the benefits of modeling the semantic dependency between hand pose and action, since intuitively action is defined by the pattern of hand motion (i.e. verb) and object in manipulation (i.e. noun). For example, [24, 31, 45, 47, 53, 52, 8] leverage the hand pose features for action recognition; while [53] further refers to the action feature for pose refinement. Furthermore, a recent work [52] demonstrates the benefit of exploiting temporal cue by respecting the temporal granularity, and proposes a framework with two cascaded blocks to respectively work on short- and long- term spans and output per-frame 3D hand pose and video action. The hierarchical structure of our framework is inspired by [52], but we have extended it to model prediction tasks, which not only covers more tasks but also enhances recognition performance (Sec. 4.3).
3D Hand and human motion prediction
Previous works predict the 2D or 3D trajectory of hand roots  [26, 25, 29, 2] or skeleton [7] from observed hand motion, with more recent works building on powerful generative deep neural networks to capture the distributions of future sequences [55, 1, 30, 4, 32, 49, 28, 20, 56]. In motion prediction, the modeling of semantic dependency also demonstrates its benefits. On one hand, [25, 9] leverage the predicted hand trajectory and its interaction with object for improved action anticipation. On the other hand, recent works also generate more realistic pose sequences by taking the past motion together with a specified action as condition, based on cVAEs [4, 32], GPT-like models [28, 20, 56] and diffusion models [49]. Our work builds a hierarchical structure for motion prediction, where consistency in both short-term motion and long-term action is explicitly ensured through the cascade of generative Transformer VAEs (Sec. 4.4). In addition, we learn prediction and recognition simultaneously, which improves both tasks by exploiting the shared temporal regularity (Sec. 4.3).
Bridging recognition and prediction
Previous studies have shown enhanced effectiveness by bridging recognition and prediction on hand (or body) poses. For example, [39, 44] build on cVAEs to model a latent space depicting the pose transition along the temporal dimension, by learning the next frame pose prediction; the learned latent space is then leveraged as a strong regulation in optimization-based refinement of estimated pose sequences. [50] models a manifold of plausible human poses based on neural distance fields; compared with Gaussian latents, the learned manifold facilitates more accurate optimization of pose refinement and improves diversity for pose generation. Rather than relying on optimization, our framework directly addresses both recognition and prediction.
Building on advances of generative models, recent works [37, 36, 28, 48, 49] learn text-guided motion generation models, where the text guidance is in the form of prescribed actions; the generated sequences can then be used for training recognition tasks. For example, PoseGPT [28] first quantizes short motion clips into latent codes by training a VQ-VAE, and then constructs a GPT-like auto-regressive model for motion generation, which learns on sequences of action and latent motion tokens. Compared with these generative models, our framework enables both pose-induced action recognition and action guided pose generation, and captures the critical semantic hierarchy between short-term pose and long-term action.
Figure 2: Overall of our framework. The cascaded $\mathbf{P}$ and $\mathbf{A}$ (denoted in the blue shadow) of G-HTT jointly model recognition and prediction, by faithfully respecting the semantic dependency and temporal granularity among pose, mid-level and action (Sec. 3.2). The two $\mathbf{D}^{P}$ in the figure are identical, but decode either $\mathbf{m}^{P}$ (gray) or $\mathbf{m}^{A}$ (white) for motion prediction.
@@ 3 Methods
The core framework, namely Generative Hierarchical Temporal Transformer (G-HTT, Fig. 2), takes as input the object in manipulation and the observed pose sequence of $T$ frames for two interacting hands (Sec. 3.1), where the hand motion and object feature respectively depict the verb and noun of the action being performed (e.g. take out a chip). G-HTT then jointly models both recognition and prediction, while following the semantic-temporal hierarchy of pose-action that captures their dependency and different temporal granularities (Sec. 3.2). In the test stage, we apply G-HTT to recognition tasks of input pose refinement and action recognition, and diverse hand motion prediction for future timestamps (Sec. 3.3). Important implementation details are given in Sec. 3.4, and a table of notations is provided in the supplementary for reference.
3.1 3D Hand Motion Representation
Given a sequence depicting two interacting hands, we construct the 3D hand representation for each frame as $\mathbf{H}=(p^{L},p^{R},v^{L},v^{R},r)$, where $p^{L},p^{R}\in\mathbb{R}^{3N}$ are the 3D coordinates for $N$ joints of each hand aligned to a predefined template [40], $v^{L},v^{R}\in\mathbb{R}^{9}$ are the relative rigid palm transformations compared with the previous frame, and $r\in\mathbb{R}^{9}$ are the relative rigid palm transformation between two hands in the current frame. We denote the rigid transformation $v^{L},v^{R},r$ by concatenating the 6D continuous representation of rotation [57] and 3D translation. More details of $\mathbf{H}$ encoding are given in the supplementary.
@@ 3.2 Joint Modeling of Recognition and Prediction with Semantic-Temporal Hierarchy
G-HTT consists of two cascaded blocks, namely the short-term pose block $\mathbf{P}$ and the long-term action block $\mathbf{A}$, to jointly model recognition and prediction while following the hierarchy of temporal and semantic granularity for pose-action. Both $\mathbf{P}$ and $\mathbf{A}$ have the same VAE structure, with their encoders and decoders respectively output for recognition and future motion prediction, but they model different semantic levels and time spans (Fig. 2).
@@
To bridge the pose and action blocks in the semantic-temporal hierarchy, we explicitly introduce a mid-level feature $\mathbf{m}$, which compresses the hand poses within a sub-second time span. $\mathbf{P}$ and $\mathbf{A}$ then respectively model the mappings between pose vs. mid-level, and mid-level vs. action. As the two blocks are cascaded, the different semantic levels can refer to each other for globally consistent recognition and prediction (Sec. 3.3). Moreover, our design enables a flexible training scenario, where $\mathbf{P}$ and $\mathbf{A}$ can be decoupled and trained separately based on their respective supervision signals and training data (Sec. 3.4).
P-Block
@@
takes a subsecond time span of $t(t<T)$ consecutive frames to model the relationship between per-frame hand pose and mid-level feature $\mathbf{m}^{P}$, without explicitly leveraging the action information. The mid-level $\mathbf{m}^{P}$ is learned to be the latent bottleneck of $\mathbf{P}$, which encodes the input $t$ consecutive frames of hand poses $\tilde{\mathbf{H}}_{1:t}$ via $\mathbf{E}^{P}$, and is decoded to hand motion ${\mathbf{H}}_{t+1:2t}$ of the future $t$ frames by $\mathbf{D}^{P}$. Meanwhile, similar to [52], the observed $\tilde{\mathbf{H}}_{1:t}$ can be refined by $\mathbf{E}^{P}$ leveraging short-term temporal regularity.
@@
In detail, $\mathbf{E}^{P}$ takes as input a sequence of $t+2$ tokens $(\tilde{\mu}^{P},\tilde{\Sigma}^{P},\tilde{\gamma_{1}},...,\tilde{\gamma_{t}})$, where $\tilde{\gamma_{i}}\in\mathbb{R}^{d}$ encodes the per-frame hand pose $\tilde{\mathbf{H}}_{i}$, and $\tilde{\mu}^{P},\tilde{\Sigma}^{P}\in\mathbb{R}^{d}$ are trainable tokens for parameterizing the distribution of $\mathbf{m}^{P}$ by aggregating over $\tilde{\mathbf{H}}_{1:t}$, similar to [37, 36]. Denoting the output sequence of $\mathbf{E}^{P}$ as $({\mu}^{P},{\Sigma}^{P},{\gamma}_{1},...,{\gamma}_{t})$, we sample $\mathbf{m}^{P}$ with re-paramterization [21] from the normal distribution $N({\mu}^{P},{\Sigma}^{P})$, and obtain ${\mathbf{H}}_{i}=\mathtt{MLP_{1}}({\gamma}_{i})$ as the refined per-frame hand poses for $i\in[1,t]$.
@@
$\mathbf{D}^{P}$ predicts the hand pose for the following $t$ frames in a parallel manner. The mid-level $\mathbf{m}^{P}$, and ${\gamma}_{1:t}$ optionally as skip-connections, are referred by $\mathbf{D}^{P}$ through cross-attention. As a parallel transformer decoder, the query input of $\mathbf{D}^{P}$ are the sinusoidal position encoding of $t$ tokens, and the output $t$ tokens are mapped by $\mathtt{MLP_{1}}$ to the future hand motion ${\mathbf{H}}_{t+1:2t}$.
@@
    $L_{comp}=\frac{1}{2t}\sum_{i=1}^{2t}||\overline{\mathbf{H}}_{i}-{\mathbf{H}}_{i}||_{1}\vspace{-1mm}$ (1)
  • The root frame trajectory loss for the predicted part:
    $L_{trj}=\frac{1}{t}\sum_{i=t+1}^{2t}\left(||\overline{s}^{L}_{i}-{s}^{L}_{i}||_{1}+||\overline{s}^{R}_{i}-{s}^{R}_{i}||_{1}\right)\vspace{-2mm}$ (2)
    where ${s}^{L}_{i},{s}^{R}_{i}\in\mathbb{R}^{9}$ are computed by accumulating $v^{L},v^{R}$ of $\mathbf{H}_{t+1:i}$ as
    $s^{L}_{i}=v^{L}_{i}\circ v^{L}_{i-1}\circ\cdots\circ v^{L}_{t+1},\ \ s^{R}_{i}=v^{R}_{i}\circ v^{R}_{i-1}\circ\cdots\circ v^{R}_{t+1}\vspace{-2mm}$ (3)
    with $\circ$ denoting composition of rigid transformation. $\overline{s}^{L}_{i},\overline{s}^{R}_{i}$ denote the GT counterpart.

### v3

# Generative Hierarchical Temporal Transformer for Hand Pose and Action Modeling

@@ Abstract
We present a novel unified framework that concurrently tackles recognition and future prediction for human hand pose and action modeling. Previous works generally provide isolated solutions for either recognition or prediction, which not only increases the complexity of integration in practical applications, but more importantly, cannot exploit the synergy of both sides and suffer suboptimal performances in their respective domains. To address this problem, we propose a generative Transformer VAE architecture to model hand pose and action, where the encoder and decoder capture recognition and prediction respectively, and their connection through the VAE bottleneck mandates the learning of consistent hand motion from the past to the future and vice versa. Furthermore, to faithfully model the semantic dependency and different temporal granularity of hand pose and action, we decompose the framework into two cascaded VAE blocks: the first and latter blocks respectively model the short-span poses and long-span action, and are connected by a mid-level feature representing a sub-second series of hand poses. This decomposition into block cascades facilitates capturing both short-term and long-term temporal regularity in pose and action modeling, and enables training two blocks separately to fully utilize datasets with annotations of different temporal granularities. We train and evaluate our framework across multiple datasets; results show that our joint modeling of recognition and prediction improves over isolated solutions, and that our semantic and temporal hierarchy facilitates long-term pose and action modeling.
Keywords:
Hand pose action modeling recognition and future prediction temporal regularity semantic and temporal hierarchy hand pose estimation hand action recognition hand motion prediction
@@ 1 Introduction
Understanding dynamic hand poses and actions is fundamental in fields such as human-robot interaction and VR/AR applications. In recent years, huge progress has been made in recognizing 3D hand poses and actions (e.g. take out a chip) from inputs such as RGB videos [66, 21, 12, 46, 58, 60]. Meanwhile, another line of research [2, 28, 29] focuses on predicting future hand motion represented as a sequence of frame-wise poses, where recent research [31] further employs generative models to achieve diverse motion prediction conditioned on a given action.
However, existing literature provides isolated solutions working on either the recognition [66, 21, 12, 46, 58, 60] or prediction [2, 28, 29, 31] side, which brings deployment complexity when integrating both sides in practical applications. Furthermore, we note that recognition and prediction tasks are naturally synergized by the temporal regularities shared through observed and future timestamps: As exemplified in Fig. 1, given the observation of hand poses reaching into the can, which indicates the action of taking out a chip, one can predict that future hand motion will describe grabbing and pulling out a chip, thus completing the action. However, isolated solutions cannot fully exploit this temporal regularity, resulting in a tendency to overfit to specific data distributions and suffer from suboptimal performances in their respective domains (Sec. 4.3).
@@
Figure 1: Jointly modeling recognition and prediction, while following the semantic dependency and temporal granularity for hand pose-action. For recognition, (1$\to$(2) moves up from short to long spans for input pose refinement and action recognition respectively. For motion prediction, two paths are available: (1$\to$(4) exploits short-term motion regularity, and (1$\to$(2$\to$(3$\to$(4) enables long-term action-guided prediction.
We present a framework with a generative Transformer VAE architecture to jointly capture both recognition and future prediction for hand pose and action modeling, therefore addressing various tasks including input 3D hand pose refinement, action recognition, and future 3D hand motion prediction. Our transformer encoder and decoder respectively produce outputs for recognition and prediction, while the VAE latent code connecting the two forces the extraction of regular and consistent hand motion and action, by predicting the future from the past and vice versa. In this way, we synergize recognition and prediction tasks and enhance the performance compared to isolated solutions (Sec. 4.3).
@@
Moreover, when it comes to pose-action modeling, extensive literature has shown the benefits of capturing the semantic dependency between the instantaneous pose and action over seconds. For example, [15, 39, 40, 52, 31] generate motion by conditioning on action to enhance realism. On the recognition side, [58, 9, 60, 51] aggregates the frame-wise hand poses estimated throughout the video to recognize the performed action. Besides modeling the semantic dependency, Wen et al. [58] further emphasize capturing the different temporal granularity of hand pose and action. Their framework, namely Hierarchical Temporal Transformer, has two cascaded encoders that capture short and long time spans respectively for effective hand pose estimation and action recognition.
@@
In addition to our joint modeling of both recognition and prediction, we are inspired by [58] to introduce block cascades, thus faithfully respecting the semantic dependency and temporal granularity of hand pose-action. To this end, we decompose our framework into two cascaded blocks that have the same generative transformer VAE structure but focus on different semantic and temporal granularities, thus naming our framework Generative Hierarchical Temporal Transformer (G-HTT): The lower pose block (P block) models hand poses over short time spans, and the upper action block (A block) models action over long time spans. A middle-level representation is further introduced to connect the two blocks: it is simultaneously the pose block VAE latent code and the action block encoder input/decoder output, and semantically encodes clip-wise motion over a subsecond span (Fig. 1).
This decomposition into block cascades offers two key advantages: First, it decouples the complex motion generation into hierarchical subtasks to respectively capture short-term and long-term temporal regularity, improving over flattened models (Sec. 4.4). Second, it brings training flexibility, as we can train the blocks separately, which not only reduces training computational cost but also allows for using datasets of different annotation granularities (Secs. 3.3 and 4.5).
@@
We train and evaluate the framework across different datasets of two-hand interactions, including H2O [24] for daily activities, Assembly101 [45] and AssemblyHands [38] for (dis-)assembling take-part toys. At test time, given a 3D hand pose sequence (e.g. per-frame estimations from the observed RGB video), we first refine it by leveraging the short-term hand motion regularity, (Fig. 1, (1)). Next, we aggregate the clip-wise motions for action recognition (Fig. 1, (1$\to$(2)). Finally, we decode the observed motions and action into a sequence of future middle-level features for motion prediction (Fig. 1, (1$\to$(2$\to$(3$\to$(4)). Evaluation results across datasets show that our framework can solve recognition problems from various camera views, and generate plausible future hand poses over time. The contribution of this paper can be summarized as follows:
@@
  • A generative Transformer VAE architecture to concurrently capture recognition and future prediction for hand pose and action modeling, which exploits the temporal regularity synergized between the past and the future, thus improving over isolated solutions.
  • A hierarchical architecture composed of two cascaded generative blocks, which models semantic dependency and temporal granularity of pose-action. This block cascade facilitates capturing both short-term and long-term temporal regularities, and further brings training flexibility.
@@
  • A comprehensive evaluation of the system on tasks such as 3D hand pose refinement, action recognition, and 3D hand motion prediction, validating the performance and design of our framework.
@@ 2 Related Works
2.0.1 Action Recognition and 3D Hand Pose Estimation
@@
Massive literature addresses perceiving hand pose and action from visual observation. For example, a series of works aims to recover the 3D hand skeleton or mesh from the visual input, where the spatial correlation within a single-frame is well exploited [66, 21, 49, 26, 36, 61], and the motion coherence along the temporal dimension is further leveraged to improve robustness under occlusion and truncation [37, 3, 10, 16, 57, 17]. Meanwhile, [5, 13, 12, 11, 46] focus on the higher semantic level, where they extract the spatial-temporal feature from the input frames to recognize the semantic hand or body action.
@@
Moreover, many works notice and exploit the benefits of modeling the semantic dependency between hand pose and action, since intuitively action is defined by the pattern of hand motion (i.e. verb) and object in manipulation (i.e. noun). For example, [27, 34, 48, 51, 60, 58, 9] leverage the hand pose features for action recognition, while Yang et al. [60] further refer to the action feature for pose refinement. Wen et al. [58] further stress capturing the respective temporal granularity of pose and action when exploiting temporal cues, and propose a framework with two cascaded blocks to respectively work on short- and long-term spans and output per-frame 3D hand pose and video action.
The hierarchical structure of our framework is inspired by [58], but we have extended it to model prediction tasks, which not only covers more tasks but also enhances recognition performance (Sec. 4.3).
2.0.2 3D Human Hand and Body Motion Prediction
Previous works predict the 2D or 3D trajectory of hand roots [29, 28, 32, 2] or skeleton [8] from the observed hand motion. Furthermore, [62, 1, 33, 4, 35, 53, 31, 22, 63] capture the distribution of future body motion with powerful generative deep neural networks. Motion prediction can also benefit from semantic dependency modeling, as achieved by taking the past motion together with a specified action as condition, based on cVAEs [4, 35], GPT-like models [31, 22, 63] and diffusion models [53]. For example, PoseGPT [31] first quantizes short motion clips into latent codes by training a VQ-VAE, and then constructs a GPT-like auto-regressive model for motion generation, which learns on sequences of action and latent motion tokens.
Our work builds a hierarchical structure for motion prediction, where consistencies in both short-term motion and long-term action are explicitly ensured through the cascade of generative Transformer VAEs (Sec. 4.4). In addition, we learn prediction and recognition simultaneously, which improves both tasks by exploiting the shared temporal regularity (Sec. 4.3).
2.0.3 Bridging Recognition and Prediction
There are previous attempts to bridge recognition and prediction, therefore benefiting recognition or prediction at the pose or action level. For example, [42, 47] learn next-frame pose prediction with cVAEs to model a latent space depicting pose transitions along the temporal dimension. Their learned latent space then serves as a strong regulation in test-time optimization for body pose estimation. [40, 39, 31, 52, 53] learn text-guided motion generation models, with the text guidance in the form of prescribed actions. The generated sequences can then be used as training data for recognition tasks. On the other hand, [14, 56, 64, 59, 7] leverage per-frame prediction for understanding high-level action, benefiting tasks such as action anticipation [14, 56] or early action detection [64, 59, 7].
In comparison, our hierarchical modeling enables capturing both recognition and prediction by modeling the semantic hierarchy between short-term pose and long-term action, which significantly boosts computational efficiency, pose/action estimation accuracy, and long-term generation fidelity. This is not considered by existing works.
@@ 3 Methods
The core framework, namely Generative Hierarchical Temporal Transformer (G-HTT,  Fig. 2), takes as input the object in manipulation and the observed pose sequence of $T$ frames for two interacting hands, where the hand motion and object feature respectively depict the verb and noun of the action being performed (e.g. take out a chip). G-HTT then jointly models both recognition and prediction, while following the semantic-temporal hierarchy of pose-action that captures their dependency and different temporal granularities (Sec. 3.1). In the test stage, we apply G-HTT to recognition tasks of input pose refinement and action recognition, and to the generation task of diverse hand motion prediction (Sec. 3.2). Important implementation details are given in Sec. 3.3. The details of our hand motion representation and a table of notations are provided in the supplementary for reference.
@@ 3.1 Joint Modeling of Recognition and Prediction with Semantic-Temporal Hierarchy
Figure 2: Overview of our framework. The cascaded $\mathbf{P}$ and $\mathbf{A}$ (shaded in blue) of G-HTT jointly model recognition and prediction, and faithfully respect the semantic dependency and temporal granularity among pose, mid-level and action (Sec. 3.1).
@@
G-HTT consists of two cascaded blocks, namely the short-term pose block $\mathbf{P}$ and the long-term action block $\mathbf{A}$, to jointly model recognition and prediction while following the hierarchy of temporal and semantic granularity for pose-action. Both $\mathbf{P}$ and $\mathbf{A}$ have the same VAE structure, with their encoders and decoders respectively outputting for recognition and future motion prediction, but $\mathbf{P}$ and $\mathbf{A}$ model different semantic levels and time spans (Fig. 2).
@@
To bridge the pose and action blocks in the semantic-temporal hierarchy, we explicitly introduce a mid-level feature $\mathbf{m}$, which represents the hand poses within a sub-second time span. $\mathbf{P}$ and $\mathbf{A}$ then respectively model the mappings between pose vs. mid-level, and mid-level vs. action. As the two blocks are cascaded, the different semantic levels can refer to each other for globally consistent recognition and prediction (Sec. 3.2). Moreover, our design enables a flexible training scenario, where $\mathbf{P}$ and $\mathbf{A}$ can be decoupled and trained separately based on their respective supervision signals and training data (Sec. 3.3).
3.1.1 P-Block
@@
takes a subsecond time span of $t$ $(t<T)$ consecutive frames to model the relationship between per-frame hand pose and mid-level feature $\mathbf{m}^{P}$, without explicitly leveraging the action information. The mid-level $\mathbf{m}^{P}$ is learned to be the latent bottleneck of $\mathbf{P}$, which encodes the input $t$ consecutive frames of hand poses, and is decoded to hand motion of the future $t$ frames. Meanwhile, similar to HTT [58], the input hand poses can be refined via the encoder $\mathbf{E}^{P}$ by leveraging the short-term temporal regularity.
@@
In detail, $\mathbf{E}^{P}$ takes as input a sequence of $t+2$ tokens $(\tilde{\mu}^{P},\tilde{\Sigma}^{P},\tilde{\mathbf{H}}_{1},...,\tilde{\mathbf{H}}_{t})$. $\tilde{\mathbf{H}}_{i}$ represents the per-frame hand pose, and $\tilde{\mu}^{P},\tilde{\Sigma}^{P}\in\mathbb{R}^{d}$ are trainable tokens for parameterizing the distribution of $\mathbf{m}^{P}$ by aggregating over $\tilde{\mathbf{H}}_{1:t}$, similar to [40, 39]. Denoting the output sequence of $({\mu}^{P},{\Sigma}^{P},\mathbf{H}_{1},...,\mathbf{H}_{t})$, we obtain $\mathbf{H}_{i}$ as the refined frame-wise hand poses, and sample $\mathbf{m}^{P}$ with re-paramterization [23] from the normal distribution $N({\mu}^{P},{\Sigma}^{P})$.
@@
$\mathbf{D}^{P}$ predicts the hand pose for the following $t$ frames in a parallel manner. The mid-level $\mathbf{m}^{P}$, with $\mathbf{H}_{1:t}$ optionally concatenated as skip-connections, are referred by $\mathbf{D}^{P}$ through cross-attention. As a parallel transformer decoder, the query input of $\mathbf{D}^{P}$ are the sinusoidal position encoding of $t$ tokens, and the output $t$ tokens are mapped into the future hand motion $\mathbf{H}_{t+1:2t}$.
@@
    $L_{comp}=\frac{1}{2t}\sum_{i=1}^{2t}||\overline{\mathbf{H}}_{i}-{\mathbf{H}}_{i}||_{1}$ (1)
  • The frame-wise root trajectory loss for the predicted part:
    $L_{trj}=\frac{1}{t}\sum_{i=t+1}^{2t}\left(||\overline{s}^{L}_{i}-{s}^{L}_{i}||_{1}+||\overline{s}^{R}_{i}-{s}^{R}_{i}||_{1}\right)$ (2)
    with $\overline{s}^{L}_{i},\overline{s}^{R}_{i}$ denote the GT counterpart.